# Comparison Principles for Stochastic Volterra equations

### v1

@@ 2. Non-negativity preserving case
¯$\overline{\mathbb{P}}[x(t)\leq y(t),t\geq 0]=\overline{\mathbb{P}}[A]\geq\limsup_{n\longrightarrow\infty}\mathbb{P}_{n}[A]=\limsup_{n\longrightarrow\infty}\mathbb{P}[X_{t}^{n}\leq Y_{t}^{n},t>0]=1.$
@@
By martingale representation theorem [34, Chapter V, Theorem 3.9], there exists, possibly on an enlargement of the probability space, a two-dimensional Brownian motion $(W^{1},W^{2})$ and a $\mathbb{R}^{2\times 2}$-valued predictable process $\alpha$ such that
@@
Then ^$\langle\widehat{\mathcal{M}}\rangle_{t}=\int_{0}^{t}\widehat{\alpha}_{s}\widehat{\alpha}_{s}^{\top}\,\mathrm{d}s=\int_{0}^{t}\alpha_{s}\alpha_{s}^{\top}\,\mathrm{d}s=\langle\mathcal{M}\rangle_{t}$. Hence $\widehat{\mathcal{M}}$ and $\mathcal{M}$ have the same law. Consequently, by (24) and [4, Lemma 2.6],
@@
where $\widehat{B}=\frac{1}{\sqrt{2}}(W^{1}+W^{2})$ is a one-dimensional Brownian motion. This completes the proof. ∎
As an application of the above theorem, we prove the existence of a non-negative continuous weak solution for a rich class of SVEs with singular initial conditions that include the rough Cox-Ingersoll-Ross process.
Theorem 2.8.
Let $(g,K,b,\sigma)$ satisfy Assumption 2.2 with $b(t,0)\geq 0$ and $\sigma(t,0)=0$ for $t\geq 0$, $K\colon(0,\infty)\longrightarrow\mathbb{R}_{+}$ is continuous, preserves non-negativity. Suppose that $g$ is such that
$g(t)=\widetilde{g}(t)+\int_{0}^{t}K(t-s)h(s)\,\mathrm{d}s$
with $0\leq h\in L_{\mathrm{loc}}^{2}(\mathbb{R}_{+})$ and $0\leq\widetilde{g}\in L_{\mathrm{loc}}^{2}(\mathbb{R}_{+})\cap C((0,\infty))$ is non-decreasing, (5) holds, and there exists $\delta\in(0,1/2)$ with (9). Then
(25) $X_{t}=g(t)+\int_{0}^{t}K(t-s)b(s,X_{s})\,\mathrm{d}s+\int_{0}^{t}K(t-s)\sigma(s,X_{s})\,\mathrm{d}B_{s}$
admits a non-negative continuous weak solution.
Proof.
The existence of a continuous weak solution $X$ follows from Theorem 2.6. Moreover, $Y_{t}=0$, $t>0$, clearly solves
$Y_{t}=\int_{0}^{t}K(t-s)(b(s,Y_{s})-b(s,0))\,\mathrm{d}s+\int_{0}^{t}K(t-s)\sigma(s,Y_{s})\,\mathrm{d}B_{s}.$
Since the pair $(0,g)$ is admissible, by Theorem 2.7 there exists a continuous weak solution $(\widehat{Y},\widehat{X})$ of the associated two-dimensional Volterra equation (23) such that $\mathcal{L}(\widehat{Y})=\mathcal{L}(Y)$, $\mathcal{L}(\widehat{X})=\mathcal{L}(X)$ and ^$\widehat{\mathbb{P}}[\widehat{Y}_{t}\leq\widehat{X}_{t},t>0]=1.$ We claim that ^$\widehat{\mathbb{P}}[\widehat{X}_{t}\geq 0,t>0]=1.$ Indeed, we have $11\geq\widehat{\mathbb{P}}[\{\widehat{Y}_{t}\geq 0,t>0\}\cup\{\widehat{Y}_{t}\leq\widehat{X}_{t},t>0\}]\geq\widehat{\mathbb{P}}[\widehat{Y}_{t}\geq 0,t>0]=$ 1. Consequently,
11 11 =^[$=\widehat{\mathbb{P}}[\{\widehat{Y}_{t}\geq 0,t>0\}\cup\{\widehat{Y}_{t}\leq\widehat{X}_{t},t>0\}]$
=^$=\widehat{\mathbb{P}}[\widehat{Y}_{t}>0,t>0]+\widehat{\mathbb{P}}[\widehat{Y}_{t}\leq\widehat{X}_{t},t>0]-\widehat{\mathbb{P}}[\{\widehat{Y}_{t}\geq 0,t>0\}\cap\{\widehat{Y}_{t}\leq\widehat{X}_{t},t>0\}]$
which yields
11 11 =^[$=\widehat{\mathbb{P}}[\{\widehat{Y}_{t}\geq 0,t>0\}\cap\{\widehat{Y}_{t}\leq\widehat{X}_{t},t>0\}]$
=^[$=\widehat{\mathbb{P}}[\{\widehat{Y}_{t}\geq 0,t>0\}\cap\{0\leq\widehat{X}_{t},t>0\}]$
^$\leq\widehat{\mathbb{P}}[0\leq\widehat{X}_{t},t>0].$
Thus, since $\mathcal{L}(\widehat{X})=\mathcal{L}(X)$, it is the desired non-negative continuous weak solution of (25). ∎
Below we illustrate this result by an application to the rough Cox-Ingersoll-Ross process.
Example 2.9 (rough Cox-Ingersoll-Ross process).
Let $\lambda,\theta,\sigma>0$, and $\alpha\in(1/2,1)$. Then for each $g$ satisfying the condition of Theorem 2.8, there exists a continuous nonnegative weak solution of
$X_{t}=g(t)+\int_{0}^{t}\frac{(t-s)^{\alpha-1}}{\Gamma(\alpha)}\lambda\left(\theta-X_{s}\right)\,\mathrm{d}s+\sigma\int_{0}^{t}\frac{(t-s)^{\alpha-1}}{\Gamma(\alpha)}\sqrt{X_{s}}\,\mathrm{d}B_{s}.$
Moreover, this process satisfies the conditions of Theorem 2.7.
Note that the non-negativity or the comparison principle holds on a newly constructed filtered probability space and subsequently does not require pathwise uniqueness for the rough Cox-Ingersoll-Ross process that is still an open problem, see also [30, 32] for related results. Moreover, we do not prove the uniqueness in law for the rough Cox-Ingersoll-Ross process constructed above. The latter would certainly follow from an extension of the affine transformation formula given in [28] applied to this setting. Since such a possible research direction goes beyond the scope of comparison principles studied in this work, we leave the details for the interested reader. However, it is relatively easy to prove the uniqueness of the time-marginals. Indeed, let $X^{g}$ be the process given by Example 2.9. Then it is easy to see that
(26) $\mathbb{E}[X^{g}_{t}]=g(t)-\lambda\int_{0}^{t}E_{\lambda}(t-s)g(s)\,\mathrm{d}s+\lambda\theta\int_{0}^{t}E_{\lambda}(s)\,\mathrm{d}s$
where $E_{\lambda}$ is the unique solution of $E_{\lambda}+\lambda K_{\alpha}\ast E_{\lambda}=K_{\alpha}$. Let $Y^{g}$ be another rough Cox-Ingersoll-Ross process with the same $g$. Since $(g,g)$ is admissible, by Theorem 2.7 we find a coupling $(\widehat{X}^{g},\widehat{Y}^{g})$ that satisfies the comparison principle, say ^$\widehat{\mathbb{P}}[\widehat{X}_{t}^{g}\leq\widehat{Y}_{t}^{g}]=1$. Hence we obtain
$W_{1}(X^{g}_{t},Y_{t}^{g})\leq\widehat{\mathbb{E}}\left[|\widehat{Y}_{t}^{g}-\widehat{X}_{t}^{g}|\right]=0$
where $W_{1}$ denotes the Wasserstein-1-distance and the last relation follows from (26). Similar arguments based on the comparison principle for Markov processes, have been used in [19, 20, 22] to prove the convergence to equilibrium including a rate of convergence. Unfortunately, due to the absence of the Markov property, an immediate extension (at least without the use of Markovian lifts) does not carry over.
@@ 3. Regular Case
where $\alpha>1$, $\beta>0$ and $x_{1},x_{2}\in\mathbb{R}$. Recall that the comparison principle holds if and only if $E_{\alpha,\beta}(-t^{\alpha})\geq 0$ for all $t\geq 0$. If $0<\beta<\alpha$ and $\alpha>1$, then $E_{\alpha,\beta}(-t^{\alpha})$ necessarily takes negative values and hence the comparison principle fails.
This example illustrates that, to obtain a comparison principle, further conditions need to be imposed. In the above example, the drift $b(x)=-x$ is non-increasing.
@@
[…]
Moreover, suppose that the ordering condition $b_{1}(x)\leq b_{2}(x)$ holds for every $x\in\mathbb{R}$ and that at least one of the functions $b_{1},b_{2}$ is monotonically non-decreasing. Let
(27) $X^{i}_{t}=g_{i}(t)+\int_{0}^{t}K(t-s)b_{i}(X^{i}_{s})\,\mathrm{d}s+\int_{0}^{t}K(t-s)\sigma(X^{i}_{s})\,\mathrm{d}B_{s}$
for $i=1,2$ be the unique strong solution. Then $\mathbb{P}[X^{1}_{t}\leq X^{2}_{t},\;t>0]=1$.
@@
Let us first note that the equations of interest have weak solutions $X^{1},X^{2}$ with sample paths in $L^{2}_{\mathrm{loc}}(\mathbb{R}_{+})$ due to [1, Theorem 1.2]. Moreover, it follows from [1, Theorem 6.1 (ii)] that $X^{1}-g_{1},X^{2}-g_{2}$ have a modification with continuous sample paths. Hence, $X^{1},X^{2}$ are continuous on $(0,\infty)$. Additionally, these solutions satisfy for any $T>0,i=1,2$, $\int_{0}^{T}\mathbb{E}[|X_{t}^{i}|^{2}]\,\mathrm{d}t<\infty$ c.f. Proposition A.1. The Yamada-Watanabe type of approach as given in [3, Proposition B.3] yields pathwise uniqueness and hence the Yamada-Watanabe-Engelbert theorem (see Kurtz [29, Theorem 1.5]) implies strong existence. Without loss of generality, we may assume that $b_{1}$ is non-decreasing. Moreover, since $K$ is non-decreasing and $g_{1},g_{2}$ are admissible,
@@
which is a continuous semimartingale. Hence we may follow the same approximation procedure as presented in [27, Chapter VI, Theorem 1.1]. That is, let $(a_{n})_{n\in\mathbb{N}_{0}}\subset(0,1]$ be a sequence with $a_{0}=1$, $a_{n}<a_{n-1}$ and $\lim_{n\longrightarrow\infty}a_{n}=0$ such that $\int_{a_{n}}^{a_{n-1}}x^{-2\eta}\,\mathrm{d}x=n.$ Then, we can construct a sequence $(\varphi_{n})_{n\in\mathbb{N}}\subset C^{2}(\mathbb{R})$ of twice continuously differentiable functions such that
@@
By using $K(0)>0$, $b_{1}(t,x)\leq b_{2}(t,x)$ and the Lipschitz continuity of $b_{1}$, we obtain
@@
It should be possible to extend the above also to the case of time-dependent coefficients. However, as the current formulation is sufficient to cover all cases of interest, we leave such an extension to the interested reader.
@@
The proof shows that the comparison principle also holds for the kernel $\widetilde{K}\coloneqq-K$ under the new assumption that $b_{1}(x)\geq b_{2}(x)$ and at least one of $b_{1},b_{2}$ is non-increasing.
The fractional kernel $K(t)=\frac{t^{\alpha-1}}{\Gamma(\alpha)}$ with $\alpha>3/2$ satisfies the assumptions stated in Theorem 3.2. The case $\alpha>1$ can be now obtained by approximation similar to Theorem 2.7. Since for such cases, the uniqueness is not guaranteed, we show that for any two weak solutions, there exists a coupling that satisfies the comparison principle.
@@
Let $(g_{i},K,b_{i},\sigma)$, $i=1,2,$ satisfy Assumption 2.2 with $b_{1}\leq b_{2}$ and $b_{1}$ or $b_{2}$ is nondecreasing. Suppose that $K\in C^{1}((0,\infty))$ with $K^{\prime}\geq 0.$ Let $g_{1},g_{2}$ be admissible such that both satisfy (9) for some $\delta\in[0,1/2)$. Then for any two weak continuous solutions $(X^{i},(\Omega^{i},\mathcal{F}^{i},\mathbb{F}^{i},\mathbb{P}^{i}),B^{i})$, $i=1,2,$ a weak continuous solution of (27), there exists a weak continuous solution $((\widehat{X}^{1},\widehat{X}^{2}),$ $(\widehat{\Omega},\widehat{\mathcal{F}},\widehat{\mathbb{F}},\widehat{\mathbb{P}}),\widehat{B})$ of the two dimensional stochastic Volterra equation
@@
such that $\mathcal{L}(\widehat{X}^{i})=\mathcal{L}(X^{i})$, $i=1,2,$ and ^$\widehat{\mathbb{P}}[\widehat{X}^{1}_{t}\leq\widehat{X}^{2}_{t},\,t>0]=1$.
@@
We choose $(b_{1,n})_{n\in\mathbb{N}}$, $(b_{2,n})_{n\in\mathbb{N}}$,$(\sigma_{n})_{n\in\mathbb{N}}$ as in the proof of Theorem 2.7 (without time dependence) and can assume that $b_{1,n}$ or $b_{2,n}$ is non-decreasing. Moreover, define $K_{n}(t)=K(t+\frac{1}{n}).$ Clearly, $K_{n}\in W^{1,2}_{\mathrm{loc}}(\mathbb{R}_{+})$ and $K_{n}\longrightarrow K$ in $L^{2}_{\mathrm{loc}}(\mathbb{R}_{+}).$ Let $(X^{n})_{n\in\mathbb{N}}$, $(Y^{n})_{n\in\mathbb{N}}$ be the family of continuous processes constructed by $(g_{1},K_{n},b_{1,n},\sigma_{n})$ and $(g_{2},K_{n},b_{2,n},\sigma_{n})$ on some fixed filtered probability space $(\Omega,\mathcal{F},\mathbb{F},\mathbb{P})$. Then, by Theorem 3.2, $\mathbb{P}[X^{n}_{t}\leq Y^{n}_{t},\;t>0].$ Using Propsition A.2, we obtain for some $p\geq 2$ suitably chosen
@@
where $\theta>0$ and $c$ is a constant that only depends $C,T,K_{1}$ and $p$. Hence, we conclude by Kolmogorov’s tightness criterion that $((X^{n}-g_{1},Y^{n}-g_{2},Z^{n}))_{n\in\mathbb{N}}$ is a tight sequence of continuous processes in $\mathbb{R}^{4}$ and so has a weakly convergent subsequence $((X^{n}-g_{1},Y^{n}-g_{2},Z^{n}))_{n\in\mathbb{N}}$ on the path space $C(\mathbb{R}_{+};\mathbb{R}^{4})$. Define the continuous mapping $\Phi\colon C(\mathbb{R}_{+};\mathbb{R}^{2})\times C(\mathbb{R}_{+};\mathbb{R}^{2})\longrightarrow C_{\delta}(\mathbb{R}_{+};\mathbb{R}^{2})\times C(\mathbb{R}_{+};\mathbb{R}^{2})$ by
@@
By the continuous mapping theorem, $(X^{n},Y^{n},Z^{n})=\Phi(X^{n}-g_{1},Y^{n}-g_{2},Z)\Longrightarrow\widetilde{\mathbb{P}}$ on $C_{\delta}(\mathbb{R}_{+};\mathbb{R}^{2})\times C(\mathbb{R}_{+};\mathbb{R}^{2})$. As in the proof of Theorem 2.7 Step 2, we can show ¯$\overline{\mathbb{P}}[x(t)\leq y(t),\;t>0]=1$ where $(x,y)$ denotes the projection to the first two coordinates of the coordinate process. Note that $C_{\delta}(\mathbb{R}_{+};\mathbb{R}^{2})\times C(\mathbb{R}_{+};\mathbb{R}^{2})$ is continuously embedded into $L^{2}_{\mathrm{loc}}(\mathbb{R}_{+};\mathbb{R}^{2})\times D(\mathbb{R}_{+};\mathbb{R}^{2})$ where we have equipped the Skorokhod space $D(\mathbb{R}_{+};\mathbb{R}^{2})$ with the usual Skorokhod topology, see e.g. [11, Section 12]. Hence by passing to a subsequence $(X^{n},Y^{n},Z^{n})\Longrightarrow(\widehat{X},\widehat{Y},\widehat{Z})$ weakly in $L^{2}_{\mathrm{loc}}(\mathbb{R}_{+};\mathbb{R}^{2})\times D(\mathbb{R}_{+};\mathbb{R}^{2})$. Using [1, Theorem 1.6] yields that $(\widehat{X},\widehat{Y},\widehat{Z})$ is a weak solution for the data $((g_{1},g_{2}),K,$ $(b_{1},b_{2}),(\sigma,\sigma)))$ and satisfies ^$\widehat{\mathbb{P}}[\widehat{X}_{t}\leq\widehat{Y}_{t},t>0]=1.$

### v2

@@ 2. Non-negativity preserving case
$\overline{\mathbb{P}}[x(t)\leq y(t),t\geq 0]=\overline{\mathbb{P}}[A]\geq\limsup_{n\to\infty}\mathbb{P}_{n}[A]=\limsup_{n\to\infty}\mathbb{P}[X_{t}^{n}\leq Y_{t}^{n},t>0]=1.$
@@
By martingale representation theorem [29, Chapter V, Theorem 3.9], there exists, possibly on an enlargement of the probability space, a two-dimensional Brownian motion $(W^{1},W^{2})$ and a $\mathbb{R}^{2\times 2}$-valued predictable process $\alpha$ such that
@@
Then $\langle\widehat{\mathcal{M}}\rangle_{t}=\int_{0}^{t}\widehat{\alpha}_{s}\widehat{\alpha}_{s}^{\top}\,\mathrm{d}s=\int_{0}^{t}\alpha_{s}\alpha_{s}^{\top}\,\mathrm{d}s=\langle\mathcal{M}\rangle_{t}$. Hence $\widehat{\mathcal{M}}$ and $\mathcal{M}$ have the same law. Consequently, by (15) and [4, Lemma 2.6],
@@
where $\widehat{B}=\frac{1}{\sqrt{2}}(W^{1}+W^{2})$ is a one-dimensional Brownian motion. This completes the proof.
Also here, let us note that the model given in Example 2.9 satisfies the conditions of Theorem 2.10, and hence fulfils the comparison principle.
@@ 3. Regular Case
where $\alpha>1$, $\beta>0$ and $x_{1},x_{2}\in\mathbb{R}$. Recall that by Example 2.6, the comparison principle holds if and only if $E_{\alpha,\beta}(-t^{\alpha})\geq 0$ for all $t\geq 0$. If $0<\beta<\alpha$ and $\alpha>1$, then $E_{\alpha,\beta}(-t^{\alpha})$ necessarily takes negative values and hence the comparison principle fails.
This example illustrates that to obtain a comparison principle, further conditions need to be imposed. In the above example, the drift b(x)=xb(x)=-x is non-increasing.
@@
Suppose $K\in W^{1,2}_{\mathrm{loc}}(\mathbb{R}_{+})$ such that $K(0)\geq 0$ and $K^{\prime}\geq 0$ [^2] By abuse of terminology, we denote by $K$ its continuous representative and by $K^{\prime}$ its weak derivative.. Let $b_{1},b_{2},\sigma\colon\mathbb{R}_{+}\times\mathbb{R}\longrightarrow\mathbb{R}$ be such that there exist constants $C>0$, $\eta\in[1/2,1]$ such that
$|b_{1}(x)-b_{1}(y)|+|b_{2}(x)-b_{2}(y)|\leq C|x-y|\quad\text{and}\quad|\sigma(x)-\sigma(y)|\leq C|x-y|^{\eta}$
hold for all $x,y\in\mathbb{R}$. Moreover, suppose that the following conditions hold:
  1. (i)
    the ordering condition $b_{1}(x)\leq b_{2}(x)$ holds for every $x\in\mathbb{R}$;
  2. (ii)
    at least one of the functions $b_{1}(\cdot),b_{2}(\cdot)$ is monotonically non-decreasing;
  3. (iii)
    Let $g_{1},g_{2}$ be given as in (4) with $\widetilde{g_{1}},\widetilde{g}_{2}\in C((0,\infty))$ such that $\widetilde{g}_{2}-\widetilde{g}_{1}\geq 0$ is non-decreasing on $(0,\infty)$, and $h_{1},h_{2}\in L^{2}_{\mathrm{loc}}(\mathbb{R}_{+})$ satisfy $h_{1}\leq h_{2}$.
Then the unique strong solutions $X^{1},X^{2}$ of (1) with $g,b$ replaced by $g_{i},b_{i}$ with $i=1,2$ satisfy $\mathbb{P}[X^{1}_{t}\leq X^{2}_{t},\;t>0]=1$.
@@
Let us first note that the equations of interest have weak solutions $X^{1},X^{2}$ with sample paths in $L^{2}_{\mathrm{loc}}(\mathbb{R}_{+})$ due to [1, Theorem 1.2]. Moreover, it follows from [1, Theorem 6.1 (ii)] that $X^{1}-g_{1},X^{2}-g_{2}$ have a modification with continuous sample paths. Hence, $X^{1},X^{2}$ are continuous on $(0,\infty)$. Additionally, these solutions satisfy for any $T>0,i=1,2$, $\int_{0}^{T}\mathbb{E}[|X_{t}^{i}|^{2}]\,\mathrm{d}t<\infty$ c.f. Proposition A.1. The Yamada-Watanabe type of approach as given in [3, Proposition B.3] yields pathwise uniqueness and hence the Yamada-Watanabe-Engelbert theorem (see Kurtz [24, Theorem 1.5]) implies strong existence.
Without loss of generality, we may assume that $b_{1}$ is non-decreasing. Moreover, since $K$ is non-decreasing by assumption (iii), we get that
@@
which is a continuous semimartingale. Hence we may follow the same approximation procedure as presented in [23, Chapter VI, Theorem 1.1]. That is, let $(a_{n})_{n\in\mathbb{N}_{0}}\subset(0,1]$ be a sequence with $a_{0}=1$, $a_{n}<a_{n-1}$ and $\lim_{n\longrightarrow\infty}a_{n}=0$ such that $\int_{a_{n}}^{a_{n-1}}x^{-2\eta}\,\mathrm{d}x=n.$ Then, we can construct a sequence $(\varphi_{n})_{n\in\mathbb{N}}\subset C^{2}(\mathbb{R})$ of twice continuously differentiable functions such that
@@
By using $K(0)>0$, $b_{1}(x)\leq b_{2}(x)$ and the Lipschitz continuity of $b_{1}$, we obtain
@@
The proof shows that the comparison principle also holds for the kernel $\widetilde{K}\coloneqq-K$ under the new assumption that $b_{1}(x)\geq b_{2}(x)$ and at least one of $b_{1}(\cdot),b_{2}(\cdot)$ is non-increasing for each $t\geq 0$.
The fractional kernel $K(t)=\frac{t^{\alpha-1}}{\Gamma(\alpha)}$ with $\alpha>3/2$ satisfies the assumptions stated in Theorem 3.2. The case $\alpha>1$ can now be obtained by approximation similar to Theorem 2.10. Since uniqueness is not guaranteed in such cases, we show the existence of a monotone coupling.
@@
[…]
@@
such that $\widehat{\mathbb{P}}[\widehat{X}^{1}_{t}\leq\widehat{X}^{2}_{t},\,t>0]=1$.
@@
We choose $(b_{1,n})_{n\in\mathbb{N}}$, $(b_{2,n})_{n\in\mathbb{N}}$$(\sigma_{n})_{n\in\mathbb{N}}$ as in the proof of Theorem 2.10 and can assume that $b_{1,n}$ or $b_{2,n}$ is non-decreasing. Moreover, define $K_{n}(t)=K(t+\frac{1}{n}).$ Clearly, $K_{n}\in W^{1,2}_{\mathrm{loc}}(\mathbb{R}_{+})$ and $K_{n}\longrightarrow K$ in $L^{2}_{\mathrm{loc}}(\mathbb{R}_{+}).$ Let $(X^{n})_{n\in\mathbb{N}}$, $(Y^{n})_{n\in\mathbb{N}}$ be the family of continuous processes constructed by $(g_{1},K_{n},b_{1,n},\sigma_{n})$ and $(g_{2},K_{n},b_{2,n},\sigma_{n})$ on some fixed filtered probability space $(\Omega,\mathcal{F},\mathbb{F},\mathbb{P})$. Then, by Theorem 3.2, $\mathbb{P}[X^{n}_{t}\leq Y^{n}_{t},\;t>0].$ Using Propsition A.2, we obtain for some $p\geq 2$ suitably chosen
@@
where $\theta>0$ and $c$ is a constant that only depends $C,T,K_{1}$ and $p$. Hence, we conclude by Kolmogorov’s tightness criterion that $((X^{n}-g_{1},Y^{n}-g_{2},Z^{n}))_{n\in\mathbb{N}}$ is a tight sequence of continuous processes in $\mathbb{R}^{4}$ and so has a weakly convergent subsequence $((X^{n}-g_{1},Y^{n}-g_{2},Z^{n}))_{n\in\mathbb{N}}$ on the path space $C(\mathbb{R}_{+};\mathbb{R}^{4})$.
Define the continuous mapping $\Phi\colon C(\mathbb{R}_{+};\mathbb{R}^{2})\times C(\mathbb{R}_{+};\mathbb{R}^{2})\longrightarrow C_{\delta}(\mathbb{R}_{+};\mathbb{R}^{2})\times C(\mathbb{R}_{+};\mathbb{R}^{2})$ by
@@
[…]
We expect that the results presented in section 3 can be extended to time-dependent coefficients. However, establishing such an extension would require extending the results in [1] to time-dependent coefficients, which is beyond the scope of this paper, and leave such an extension to the interested reader.